# A Statistical Framework of Watermarks for Large Language Models: Pivot, Detection Efficiency and Optimal Rules

### v1

@@ 3 Application to the Gumbel-max Watermark
Uniqueness of the Gumbel-max decoder.
@@
Suppose $V$ satisfies both (16) and (17). Then ${\mathcal{S}}^{V}$ is unbiased if and only if
@@ 3.1 Main Results
Optimal score function.
@@
We write $Y^{\mathrm{gum}}_{t}\equiv Y^{\mathrm{gum}}({w}_{t},\zeta_{t}):=U_{t,{w}_{t}}$. The distribution of this pivot under $H_{1}$ is a mixture of Beta distributions, which is a classic result in the literature.
@@
Figure 2: Illustration of the distribution of $Y^{\mathrm{gum}}_{t}$. We plot the CDF $F_{1,\bm{P}_{t}}$ and the PDF $f_{\bm{P}_{t}}$ by assuming $P_{t,{w}}\propto({w}+b)^{-a},{w}\in{{\mathcal{W}}}$ with $|{{\mathcal{W}}}|=15$ and different choices of $(a,b)$. The resulting distribution is denoted by $\mathrm{Pow}(a,b)$. $F_{1,\bm{P}_{t}}$ deviates a lot from the uniform distribution.
The fact that this distribution is stochastically larger than $U(0,1)$ can be gleaned from the following inequality:
$\sum_{{w}\in{{\mathcal{W}}}}P_{t,{w}}r^{1/P_{t,{w}}}=r\sum_{{w}\in{{\mathcal{W}}}}P_{t,{w}}r^{1/P_{t,{w}}-1}\leq r\sum_{{w}\in{{\mathcal{W}}}}P_{t,{w}}=r.$
Also, see Figure 2 for an illustration: there is a higher accumulation of probability mass around one under $H_{1}$ compared to the distribution observed under $H_{0}$. From this observation, any score function $h$ aware of the distributional differences could be used for detecting the Gumbel-max watermarks. For example, [1] uses $h_{\mathrm{ars}}(r)=-\log(1-r)$, while [36, 22] propose $h_{\mathrm{log}}(r)=\log r$. All of these score functions are increasing functions that assign large values to points near one. While all these choices control the Type I error, a crucial question is on what basis we compare score functions and which is the optimal one.
@@
This function gives the optimal ${\mathcal{P}}_{\Delta}$-efficiency rates in the sense that
$R_{{\mathcal{P}}_{\Delta}}({h_{\mathrm{gum},\Delta}^{\star}})\geq R_{{\mathcal{P}}_{\Delta}}(h)$
for any measurable function $h$. Moreover, $R_{{\mathcal{P}}_{\Delta}}({h_{\mathrm{gum},\Delta}^{\star}})$ is attained at the following least-favorable NTP distribution in ${\mathcal{P}}_{\Delta}$:
@@
To clear off any confusion, the second conclusion amounts to saying that $({h_{\mathrm{gum},\Delta}^{\star}},\bm{P}_{\Delta}^{\star})$ is a saddle point solution to the minimax problem (14). Explicitly, $\bm{P}$ maximizes $L({h_{\mathrm{gum},\Delta}^{\star}},\bm{P})$ in (14) if and only if $\bm{P}$ can be obtained by permuting the coordinates of $\bm{P}_{\Delta}^{\star}$.
@@
As an interesting fact, the least-favorable distribution $\bm{P}_{\Delta}^{\star}$ and its permuted counterparts form all the vertices of ${\mathcal{P}}_{\Delta}$ (see Lemma 3.4 in Section 3.2). Moreover, the vertices are the closest NTP distributions in ${\mathcal{P}}_{\Delta}$ to singular distributions. Roughly speaking, the closer to singular distributions, the more difficult it is to test apart the two hypotheses in (9). To see this, note that the testing problem becomes most difficult when the NTP distribution is singular, which makes the null and alternative distributions of the pivot indistinguishable.
Comparison with other rules.
@@
For any non-decreasing function $h$, the functional [^12] We consider a slightly different version of the MGF by setting $\theta=1$ and taking as input $\bm{P}$ in (11).
$\bm{P}\mapsto\phi_{h}(\bm{P}):=\operatorname{\mathbb{E}}_{1,\bm{P}}{\mathrm{e}}^{-h(Y^{\mathrm{gum}})}$
is convex in $\bm{P}\in{\mathrm{Simp}}({{\mathcal{W}}})$.
This convexity lemma is the key to the proof of Theorem 3.2 and is a contribution of independent interest to future work on LLM watermarks. Roughly speaking, it helps reduce the “max” part in the minimax problem (14) to the problem of identifying vertices of the distribution class.
Technically speaking, the role of the convexity lemma in the proof of Theorem 3.2 is through the following lemma.
@@
A direct consequence of Lemma 3.3 is that
$R_{{\mathcal{P}}_{\Delta}}(h)=-\inf_{\theta\geq 0}L(\theta h,\bm{P}_{\Delta}^{\star}).$
@@
By Lemma 3.1, the CDF of $Y^{\mathrm{gum}}$ with NTP distribution $\bm{P}$ takes the form
$F_{1,\bm{P}}(r)=\sum_{{w}\in{{\mathcal{W}}}}P_{{w}}r^{1/P_{{w}}}~{}~{}\text{for $r\in[0,1]$}.$
This function relates to $\phi_{h}(\bm{P})$ since
$\phi_{h}(\bm{P})=\operatorname{\mathbb{E}}_{1,\bm{P}}{\mathrm{e}}^{-h(Y^{\mathrm{gum}})}=\int_{0}^{1}{\mathrm{e}}^{-h(r)}F_{1,\bm{P}}({\mathrm{d}}r)$
@@
Our proof is based on a fundamental principle in convex analysis: the supremum of a convex function over a compact convex set in an Euclidean space is necessarily attained at an extreme point of the convex set [28].
Fix a non-decreasing function $h$. The mapping $\bm{P}\mapsto\phi_{h}(\bm{P})$ is shown to be convex in $\bm{P}$ by Lemma 3.3. Additionally, ${\mathcal{P}}_{\Delta}$ is convex. Thus the supremum of the convex function $\bm{P}\mapsto\phi_{h}(\bm{P})$ must be necessarily attained at an extreme point of the convex constraint set ${\mathcal{P}}_{\Delta}$. By Lemma 3.4, the extreme points of the set ${\mathcal{P}}_{\Delta}$ equal to all possible permutations of $\bm{P}_{\Delta}^{\star}$.
@@
$\geq\min_{h}\left(\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{gum}})+\log\phi_{h}(\bm{P}_{\Delta}^{\star})\right)$
$=-D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}_{\Delta}^{\star}}),$
where second equality is due to the Donsker–Varadhan representation [20] and $D_{\mathrm{KL}}$ denotes the Kullback–Leibler divergence. Note that Lemma 3.3 can not be applied here because the minimization function $h$ is not necessarily non-decreasing.
@@
$=\operatorname{\mathbb{E}}_{0}{h_{\mathrm{gum},\Delta}^{\star}}(Y^{\mathrm{gum}})+\log\phi_{{h_{\mathrm{gum},\Delta}^{\star}}}(\bm{P}_{\Delta}^{\star})$
$=-D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}_{\Delta}^{\star}}),$
where the second equality follows from Lemma 3.3 and the third equality is ensured by, again, the Donsker–Varadhan representation.
@@ 4 Application to the Inverse Transform Watermark
where $\zeta_{t}=(\pi_{t},U_{t})$ with $U_{t}\sim U(0,1)$ and $\pi_{t}$ being sampled uniformly at random from all permutations on ${{\mathcal{W}}}$. Following the strategy in [47, 36], $(\pi_{t},U_{t})$’s are jointly independent across the token sequence. For a given permutation $\pi$, $F(\cdot,\pi)$ above denotes the CDF under permutation:
$F(x;\pi)=\sum_{{w}^{\prime}\in{{\mathcal{W}}}}P_{{w}^{\prime}}\cdot{\mathbf{1}}_{\{\pi({{w}^{\prime}})\leq x\}}.$
By construction, the inverse transform watermark is unbiased for sampling from the NTP distribution $\bm{P}$. To see the unbiasedness directly, for any token $w$, note that ${\mathcal{S}}^{\mathrm{inv}}(\bm{P},\zeta)={w}$ if and only if
@@ 4.1 Main Results
The first part of Lemma 4.1 in particular shows that $Y_{t}^{{\mathrm{dif}}}$ is a pivotal statistic under $H_{0}$. The second part of Lemma 4.1 provides the explicit CDF of $Y_{t}^{{\mathrm{dif}}}$ under $H_{1}$ conditional on $\bm{P}_{t}$.
@@
Optimal score function.
@@
Then,
$\lim_{M\to\infty}\overline{R}_{\Delta}\left([{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}\right)=\infty.$ (26)
@@
Finally, the truncation in (26) is mainly for technical reasons, stemming from a proof artifact.
@@
Fix $\Delta\in(0,1]$. To prove Theorem 4.2, our goal is to establish the limit:
$\lim_{M\to\infty}\overline{R}_{\Delta}\left([{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}\right)=\infty.$
Our strategy involves establishing an effective lower bound for $\overline{R}_{\Delta}(h)$, as detailed in Lemma 4.2. This involves introducing a function $F_{{\mathrm{dif}},\Delta}$ for each $\Delta\in[0,1]$, defined as:
$F_{{\mathrm{dif}},\Delta}(r)=1-\max\left\{1-\frac{r}{1-\Delta},0\right\}^{2}~{}~{}\text{for any~{}~{}$r\in[0,1]$}.$
By Theorem 4.1, this corresponds to the limiting CDF of $Y^{\mathrm{dif}}$ given $\bm{P}_{t}$, assuming that $\bm{P}_{t}\in\overline{{\mathcal{P}}}_{\Delta}$.
@@
Nonetheless, after truncating ${h_{\mathrm{dif},\Delta}^{\star}}$ to $h_{M}=[{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}$, the function $h_{M}$ is non-decreasing, Lipschitz-continuous, and uniformly bounded on ${{\mathbb{R}}}$ for any $M<\infty$. This allows us to use Lemma 4.2 to conclude that for every $M<\infty$:
$\overline{R}_{\Delta}(h_{M})\geq G(h_{M}).$
Applying the limit $M\to\infty$ on both sides, and leveraging Lebesgue’s dominated convergence theorem along with Fatou’s lemma, we deduce:
@@
We are interested in lower bounding $\overline{R}_{\Delta}(h)$. Recall that its definition is given by:
$\begin{split}\overline{R}_{\Delta}(h)&=\liminf R_{\overline{{\mathcal{P}}}_{\Delta}}(h)=-\limsup\inf_{\theta\geq 0}\sup_{\bm{P}\in\overline{{\mathcal{P}}}_{\Delta}}\left[\theta\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{dif}})+\log\operatorname{\mathbb{E}}_{1,\bm{P}}[{\mathrm{e}}^{-\theta h(Y^{\mathrm{dif}})}]\right].\end{split}$ (28)
Our lower bound strategy relies on first connecting $\overline{R}_{\Delta}(h)$ to an intermediate quantity:
@@
For every function $h$ Lipschitz-continuous on $[0,1]$, we have:
@@
Lemma 4.3 implies that maximizing $\overline{R}_{\Delta}(h)$ is related to the following minimax optimization:
$\inf_{h}\sup_{\Delta^{\prime}\geq\Delta}\overline{L}(h,\Delta^{\prime}).$
This minimax formulation is closely related to our original minimax formulation (14), yet it replaces the $|{{\mathcal{W}}}|$-dimensional variable $\bm{P}$ with a scalar $\Delta^{\prime}$ after taking asymptotics (Theorem 4.1).
@@
After we substitute this inequality into (28), we obtain a lower bound:
@@
The caveat is that the above, uniform limit applies only to functions $h$ that are Lipschitz-continuous and uniformly bounded on ${{\mathbb{R}}}$, as required in Lemma 4.5.
@@
For any non-increasing and Lipschitz-continuous $h$ on $[0,1]$:
@@
Using integration by parts, we first obtain the following equation:

### v2

@@ 3 Application to the Gumbel-max Watermark
Uniqueness of the Gumbel-max decoder
@@
Suppose $V$ satisfies both (16) and (17). ${\mathcal{S}}^{V}$ is unbiased if and only if
@@ 3.1 Main Results
Optimal score function
@@
We write $Y^{\mathrm{gum}}_{t}\equiv Y^{\mathrm{gum}}({w}_{t},\zeta_{t}):=U_{t,{w}_{t}}$. This choice of pivot can be seen as an approximation of the most “optimistic” log-likelihood ratio used in an independent hypothesis test for watermark detection (see Appendix B.1 for details). The distribution of this pivot under $H_{1}$ is a mixture of Beta distributions, which is a classic result in the literature.
@@
Figure 2: Illustration of the distribution of $Y^{\mathrm{gum}}_{t}$. We plot the CDF $F_{1,\bm{P}_{t}}$ and the PDF $f_{\bm{P}_{t}}$ by assuming $P_{t,{w}}\propto({w}+b)^{-a},{w}\in{{\mathcal{W}}}$ with $|{{\mathcal{W}}}|=15$ and different choices of $(a,b)$. The resulting distribution is denoted by $\mathrm{Pow}(a,b)$. $F_{1,\bm{P}_{t}}$ deviates from the uniform distribution.
The fact that this distribution is stochastically larger than $U(0,1)$ can be gleaned from the following inequality: $\sum_{{w}\in{{\mathcal{W}}}}P_{t,{w}}r^{1/P_{t,{w}}}=r\sum_{{w}\in{{\mathcal{W}}}}P_{t,{w}}r^{1/P_{t,{w}}-1}\leq r\sum_{{w}\in{{\mathcal{W}}}}P_{t,{w}}=r.$ Also, see Figure 2 for an illustration: there is a higher accumulation of probability mass around one under $H_{1}$ compared to the distribution observed under $H_{0}$. From this observation, any score function $h$ aware of the distributional differences could be used for detecting the Gumbel-max watermarks. For example, [1] uses $h_{\mathrm{ars}}(r)=-\log(1-r)$, while [38, 22] propose $h_{\mathrm{log}}(r)=\log r$. All of these score functions are increasing functions that assign large values to points near one. While all these choices control the Type I error, a crucial question is on what basis we compare score functions and which is the optimal one.
@@
This function gives the optimal ${\mathcal{P}}_{\Delta}$-efficiency rates in the sense that $R_{{\mathcal{P}}_{\Delta}}({h_{\mathrm{gum},\Delta}^{\star}})\geq R_{{\mathcal{P}}_{\Delta}}(h)$ for any measurable function $h$. Moreover, $R_{{\mathcal{P}}_{\Delta}}({h_{\mathrm{gum},\Delta}^{\star}})$ is attained at the following least-favorable NTP distribution in ${\mathcal{P}}_{\Delta}$:
@@
To clear up any confusion, the second conclusion amounts to saying that $({h_{\mathrm{gum},\Delta}^{\star}},\bm{P}_{\Delta}^{\star})$ is a saddle point solution to the minimax problem (14). Explicitly, $\bm{P}$ maximizes $L({h_{\mathrm{gum},\Delta}^{\star}},\bm{P})$ in (14) if and only if $\bm{P}$ results from rearranging $\bm{P}_{\Delta}^{\star}$’s coordinates.
@@
As an interesting fact, the least-favorable distribution $\bm{P}_{\Delta}^{\star}$ and its permuted counterparts form all the vertices of ${\mathcal{P}}_{\Delta}$ (see Lemma 3.4 in Section 3.2). Moreover, the vertices are the closest NTP distributions in ${\mathcal{P}}_{\Delta}$ to singular distributions. Roughly speaking, the closer to singular distributions, the more difficult it is to tease apart the two hypotheses in (9). To see this, note that the testing problem becomes most difficult when the NTP distribution is singular, which makes the null and alternative distributions of the pivot indistinguishable.
Comparison with other rules
@@
For any non-decreasing function $h$, the functional [^15] We consider a slightly different version of the MGF by setting $\theta=1$ and taking as input $\bm{P}$ in (11). $\bm{P}\mapsto\phi_{h}(\bm{P}):=\operatorname{\mathbb{E}}_{1,\bm{P}}{\mathrm{e}}^{-h(Y^{\mathrm{gum}})}$ is convex in $\bm{P}\in{\mathrm{Simp}}({{\mathcal{W}}})$.
This convexity lemma is the key to the proof of Theorem 3.2 and is a contribution of independent interest to future work on LLM watermarks. Roughly speaking, it helps reduce the “max” part in the minimax problem (14) to the problem of identifying vertices of the distribution class. Technically speaking, the role of the convexity lemma in the proof of Theorem 3.2 is through the following lemma.
@@
A direct consequence of Lemma 3.3 is that $R_{{\mathcal{P}}_{\Delta}}(h)=-\inf_{\theta\geq 0}L(\theta h,\bm{P}_{\Delta}^{\star}).$
@@
By Lemma 3.1, the CDF of $Y^{\mathrm{gum}}$ with NTP distribution $\bm{P}$ takes the form: $F_{1,\bm{P}}(r)=\sum_{{w}\in{{\mathcal{W}}}}P_{{w}}r^{1/P_{{w}}}$ for $r\in[0,1]$. This function relates to $\phi_{h}(\bm{P})$ since
$\phi_{h}(\bm{P})=\operatorname{\mathbb{E}}_{1,\bm{P}}{\mathrm{e}}^{-h(Y^{\mathrm{gum}})}=\int_{0}^{1}{\mathrm{e}}^{-h(r)}F_{1,\bm{P}}({\mathrm{d}}r).$
@@
Our proof is based on a fundamental principle in convex analysis: the supremum of a convex function over a compact convex set in a Euclidean space is necessarily attained at an extreme point of the convex set [29].
Fix a non-decreasing function $h$. The mapping $\bm{P}\mapsto\phi_{h}(\bm{P})$ is shown to be convex in $\bm{P}$ by Lemma 3.3. Additionally, ${\mathcal{P}}_{\Delta}$ is convex. Thus, the supremum of the convex function $\bm{P}\mapsto\phi_{h}(\bm{P})$ must be necessarily attained at an extreme point of the convex constraint set ${\mathcal{P}}_{\Delta}$. By Lemma 3.4, any extreme point of the set ${\mathcal{P}}_{\Delta}$ must be a permuted $\bm{P}_{\Delta}^{\star}$.
@@
$\geq\min_{h}\left(\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{gum}})+\log\phi_{h}(\bm{P}_{\Delta}^{\star})\right)=-D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}_{\Delta}^{\star}}),$
where the final equality is due to the Donsker–Varadhan representation [20] and $D_{\mathrm{KL}}$ denotes the Kullback–Leibler divergence. Note that Lemma 3.3 can not be applied here because the minimization function $h$ is not necessarily non-decreasing.
@@
$=\operatorname{\mathbb{E}}_{0}{h_{\mathrm{gum},\Delta}^{\star}}(Y^{\mathrm{gum}})+\log\phi_{{h_{\mathrm{gum},\Delta}^{\star}}}(\bm{P}_{\Delta}^{\star})=-D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}_{\Delta}^{\star}}),$
where the final equality follows from Lemma 3.3 and the third equality is ensured by, again, the Donsker–Varadhan representation.
@@ 4 Application to the Inverse Transform Watermark
where $\zeta_{t}=(\pi_{t},U_{t})$ with $U_{t}\sim U(0,1)$ and $\pi_{t}$ being sampled uniformly at random from all permutations on ${{\mathcal{W}}}$. Following the strategy in [49, 38], $(\pi_{t},U_{t})$’s are jointly independent across the token sequence. For a given permutation $\pi$, $F(\cdot,\pi)$ above denotes the CDF under permutation: $F(x;\pi)=\sum_{{w}^{\prime}\in{{\mathcal{W}}}}P_{{w}^{\prime}}\cdot{\mathbf{1}}_{\{\pi({{w}^{\prime}})\leq x\}}.$ By construction, the inverse transform watermark is unbiased for sampling from the NTP distribution $\bm{P}$. To see the unbiasedness directly, for any token $w$, note that ${\mathcal{S}}^{\mathrm{inv}}(\bm{P},\zeta)={w}$ if and only if
@@ 4.1 Main Results
In Lemma 4.1, the first part in particular shows that $Y_{t}^{{\mathrm{dif}}}$ is a pivotal statistic under $H_{0}$. The second part provides the explicit CDF of $Y_{t}^{{\mathrm{dif}}}$ under $H_{1}$ conditional on $\bm{P}_{t}$.
@@
Optimal score function
@@
Then, $\lim_{M\to\infty}\overline{R}_{\Delta}\left([{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}\right)=\infty.$
@@
Finally, the truncation in $[{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}$ is mainly for technical reasons, stemming from a proof artifact.
@@
Fix $\Delta\in(0,1]$. To prove Theorem 4.2, our goal is to establish the limit: $\lim_{M\to\infty}\overline{R}_{\Delta}\left([{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}\right)=\infty.$ Our strategy involves establishing an effective lower bound for $\overline{R}_{\Delta}(h)$, as detailed in Lemma 4.2. This involves introducing a function $F_{{\mathrm{dif}},\Delta}$ for each $\Delta\in[0,1]$, defined as: $F_{{\mathrm{dif}},\Delta}(r)=1-\max\left\{1-\frac{r}{1-\Delta},0\right\}^{2}$ for any $r\in[0,1]$. By Theorem 4.1, this corresponds to the limiting CDF of $Y^{\mathrm{dif}}$ given $\bm{P}_{t}$, assuming that $\bm{P}_{t}\in\overline{{\mathcal{P}}}_{\Delta}$.
@@
Nonetheless, after truncating ${h_{\mathrm{dif},\Delta}^{\star}}$ to $h_{M}=[{h_{\mathrm{dif},\Delta}^{\star}}]_{[-M,M]}$, the function $h_{M}$ is non-decreasing, Lipschitz-continuous, and uniformly bounded on ${{\mathbb{R}}}$ for any $M<\infty$. This allows us to use Lemma 4.2 to conclude that for every $M<\infty$: $\overline{R}_{\Delta}(h_{M})\geq G(h_{M}).$ Applying the limit $M\to\infty$ on both sides, and leveraging Lebesgue’s dominated convergence theorem along with Fatou’s lemma, we deduce:
@@
We are interested in lower bounding $\overline{R}_{\Delta}(h)=\liminf R_{\overline{{\mathcal{P}}}_{\Delta}}(h)$. Recall that its definition is given by:
$\begin{split}\overline{R}_{\Delta}(h)=-\limsup\inf_{\theta\geq 0}\sup_{\bm{P}\in\overline{{\mathcal{P}}}_{\Delta}}\left[\theta\operatorname{\mathbb{E}}_{0}h(Y^{\mathrm{dif}})+\log\operatorname{\mathbb{E}}_{1,\bm{P}}[{\mathrm{e}}^{-\theta h(Y^{\mathrm{dif}})}]\right].\end{split}$ (27)
Our lower bound strategy relies on first connecting $\overline{R}_{\Delta}(h)$ to an intermediate quantity that
@@
For every function $h$ Lipschitz-continuous on $[0,1]$, we have
@@
Lemma 4.3 implies that maximizing $\overline{R}_{\Delta}(h)$ is related to the minimax optimization: $\inf_{h}\sup_{\Delta^{\prime}\geq\Delta}\overline{L}(h,\Delta^{\prime}).$ This minimax formulation is closely related to our original minimax formulation (14), yet it replaces the $|{{\mathcal{W}}}|$-dimensional variable $\bm{P}$ with a scalar $\Delta^{\prime}$ after taking asymptotics (Theorem 4.1).
@@
After we substitute this inequality into (27), we obtain a lower bound
@@
The caveat is that the above uniform limit applies only to functions $h$ that are Lipschitz-continuous and uniformly bounded on ${{\mathbb{R}}}$, as required in Lemma 4.5.
@@
For any non-increasing and Lipschitz-continuous $h$ on $[0,1]$,
@@
Using integration by parts, we first obtain the equation: